Reports 1-1 of 1 Clear search Modify search
takahiro.yamamoto - 20:09 Monday 02 May 2022 (20639) Print this report
Comment to ITMX F3 GAS signals became glitchy from yesterday (19212)

[Ushiba, YamaT]


- Similar issue on ITMY ACC_H2 was reported as klog#20636.
- Glitches on F3 GAS were removed by replacing the ADC board, the ADC adapter board, and the ADC internal cable.
- Glitch problem was solved.

- On the other hand, ADC timing error is occurred with new ADC.
- It shouldn't affect ITMY local measurement.
- Timing error probably affects measurements with channels on other real-time computer.
- This will be discussed in klog#20642.


Differences from klog#20636
Replaced info.
- ADC: S1402718 -> S2113340.

Time series of ADC noise on GAS channels are shown in Fig.1.
Behavior of F3_GAS was drastically improved and noise level of other channels were also slightly imporved.

According to spectra show in Fig.2, noise level became roughly two times smaller than ones before replacement.
(6e-3 ct/rHz => 3e-3 ct/rHz above ~50Hz)

Images attached to this comment
Search Help